A cyclist enters a curve of 30 m radius at a speed of 12 m s-1. He applies the
brakes and decreases his speed at a constant rate of 0.5 m s-2. Calculate the
cyclist’s centripetal (radial) and tangential accelerations when he is travelling at a speed of 10 m s-1.
1
Expert's answer
2019-03-13T11:53:56-0400

Solution:

We can be found the radial acceleration from the formula:


"a_r=v^2\/r"

So, we get

"a_r=10^2\/30=3.3 m\/s^2"

A cyclist reduces speed at a speed of 0.5 ms-2 This means that at any time or speed its tangential acceleration is 0.5. Therefore, when he travels at a rate of 10 ms-1, his tangential acceleration is -0.5 ms-2 because he slows down.


"a_t = -0.5 m\/s^2"

Answer: 3.3 ms-2, -0.5 ms-2